# Global well-posedness for the generalized intermediate NLS with a nonvanishing condition at infinity

### v1

@@ 1. Introduction
We study the intermediate nonlinear Schrödinger equation,
@@
The equation (INLS) was introduced by Pelinovsky in [P] to describe the long-term dynamics of the envelope of weakly nonlinear internal waves in a stratified fluid of finite depth.
@@
Thus, as shown in [CFL, Theorem 1.3], when $\delta\to\infty$, (INLS) reduces to the defocusing Calogero-Moser derivative nonlinear Schrödinger equation
@@
originally derived on the real line in [P, Matsuno2, GL], and extended to the periodic setting in [Matsuno3, Badreddine1, Badreddine2]. The equation (CM-defoc) has been deeply studied in recent years, where global well-posedness has been derived in the Hardy space $L^{2}_{+}(\mathbb{R})$ in [killip] and in $L^{2}_{+}(\mathbb{T})$ in [Badreddine1]. The focusing variant
@@
has also been shown to be globally well-posed in $L^{2}_{+}(\mathbb{R})$ [GL, killip] for initial data $\|u(0)\|_{L^{2}}^{2}<2\pi$ and in $L^{2}_{+}(\mathbb{T})$ for small initial data [Badreddine1], with the bound $2\pi$ known to be optimal thanks to the work of [Hogan-Kowalski, kim-kim-kwon, kim-Kwon, jJeong-Kim]. Recently, Chapouto–Forlano–Laurens [CFL] proved that the well-posedness theory for (CM-foc) extends beyond the Hardy-space setting. They established local well-posedness in $H^{s}(\mathbb{R})$ for all $s>1/4$, and global well-posedness for initial data with small $L^{2}$-norm. Solitons for (CM-foc) were derived in [GL, Matsuno2, Matsuno4, Matsuno9, Matsuno5] on the real line, and they display turbulent behavior [GL]. On the torus, the set of special solutions, including traveling waves, is much wider than on $\mathbb{R}$ [Badreddine2, Matsuno2, Matsuno9, Matsuno5]. We also mention that the zero dispersion limit on $\mathbb{R}$ was studied in [Matsuno3, Badreddine3], and the inverse scattering theory for the Calogero-Moser model in the defocusing [Matsuno7] and focusing [Frank-Read, chen2] cases has also been investigated.
@@
On the other hand, in the shallow-water limit $\delta\to 0$, after a suitable rescaling, (INLS) formally converges to the defocusing nonlinear Schrödinger equation
(NLS) $i\partial_{t}u-\partial_{x}^{2}u=-u|u|^{2}.$
@@
This equation and the focusing variant are famous integrable equations solved by the method of inverse scattering transformation [ablowitz, shabat1972exact2, faddeev, Ablowitz2, grebert-Kappeler]. Recently, Harrop-Griffiths–Killip–Vişan [HGKV] proved that (NLS) and the focusing variant are globally well-posed in $H^{s}(\mathbb{R})$ for any s>1/2s>-1/2. This result is sharp in the sense that the norm inflation occurs in $H^{s}(\mathbb{R})$ for $s\leq-1/2$ (see, e.g., [CCT, Kishimoto]). See, e.g., [Sulem, HGKV] for the development of well-posedness of nonlinear Schrödinger equations.
@@
A notable property of (INLS) is that it admits dark solitons [P, PG, Matsuno1]: for any $\rho>0$ and integer $N\geq 1$, there exists an $N$-soliton solution $Q$ which satisfies (INLS) on the whole time interval $\mathbb{R}$ and a nonvanishing boundary condition $|Q(x,t)|\to\rho$ as $|x|\to\infty$ for all $t\in\mathbb{R}$. The existence of dark solitons naturally suggests that the appropriate functional framework for analyzing such solutions to (INLS) should involve nonvanishing boundary conditions. Accordingly, for $\rho>0$ we introduce the following subset $\mathcal{Z}^{k}_{\rho}$ of Zhidkov space:
(1.2) ${\mathcal{Z}^{k}_{\rho}}:=\{u\in L^{\infty}(\mathbb{R})\colon\partial_{x}u\in H^{k-1}(\mathbb{R}),~|u|^{2}-\rho^{2}\in L^{2}(\mathbb{R})\}.$
@@
We will use the spaces $\mathcal{Z}^{k}_{\rho}$ only for positive integers $k$. Recall that this is not a vector space but a complete metric space with the following distance function:
(1.3) $d^{k}_{\rho}(f,g):=\|\partial_{x}(f-g)\|_{H^{k-1}}+\|f-g\|_{L^{\infty}}+\big\||f|^{2}-|g|^{2}\big\|_{L^{2}}.$
To measure the size of functions in $\mathcal{Z}^{k}_{\rho}$, we introduce the functional
(1.4) $\mathcal{E}^{k}_{\rho}(u):=\|\partial_{x}u\|_{H^{k-1}}^{2}+\|u\|_{L^{\infty}}^{2}+\big\||u|^{2}-\rho^{2}\big\|_{L^{2}}.$
@@
Although the existence of dark solitons for (INLS) has been extensively studied in the literature [P, PG, Matsuno1, Matsuno2], and their asymptotic behavior has been investigated in [Akahori], the problem of local well-posedness in a functional setting appropriate for these solitons remained open. In [DeMoura], local well-posedness222 Actually, he proved the local well-posedness of a family of nonlocal nonlinear Schrödinger equation [pelinovsky-Grimshaw] that includes also (INLS). was established in the Sobolev space $H^{s}(\mathbb{R})$ for $s\geq 1$. Later, Barros–de Moura–Santos have shown in [B-deM-S] the local well-posedness of (INLS) for small initial data in the Besov space $B_{2}^{1/2,1}(\mathbb{R})$. More recently, Chapouto–Forlano–Laurens have proven in [CFL] that (INLS) is globally well-posed in $H^{s}(\mathbb{R})$ with $s>1/4$ for small $L^{2}$ initial data. However, this setting does not accommodate the nonvanishing boundary conditions intrinsic to dark solitons. The aim of the present work is therefore to address this gap by proving that (INLS) is globally well-posed in the space $\mathcal{Z}^{2}_{\rho}$.
@@
Our first main result establishes local well-posedness in $\mathcal{Z}^{2}_{\rho}$ for (gINLS).
Theorem 1.1 (Local well-posedness in $\mathcal{Z}^{2}_{\rho}$).
Let $\delta>0$, $\alpha,\beta\in\mathbb{R}$ be arbitrary. For $M_{0}>0$, there exists $T_{*}\in(0,1]$ satisfying $T_{*}\geq c_{\alpha,\beta}(1+\delta^{-1})^{-1}(1+M_{0}^{4})^{-1}$ such that the following holds.
For any $\rho>0$ and $\phi\in\mathcal{Z}^{2}_{\rho}$ with $\mathcal{E}^{2}_{\rho}(\phi)\leq M_{0}^{2}$, there exists a unique solution $u\in C([0,T_{*}];\mathcal{Z}^{2}_{\rho})$ to (gINLS) on $0,T_{*}]$ with initial condition $u(0)=\phi$. Moreover, the mapping $\phi\mapsto u$ from $\{f\in\mathcal{Z}^{2}_{\rho}:\mathcal{E}^{2}_{\rho}(f)\leq M_{0}^{2}\}$ into $C([0,T_{*}];\mathcal{Z}^{2}_{\rho})$ is continuous. Furthermore, if in addition $\phi\in\mathcal{Z}^{k}_{\rho}$ for some integer $k\geq 3$, then $u\in C([0,T_{*}];\mathcal{Z}^{k}_{\rho})$.
@@
maps a solution $u$ of (INLS) to another solution of (INLS), but with rescaled depth $\lambda\delta$. This formally indicates that $L^{2}(\mathbb{R})$ is the scaling-critical space. Consequently, one may expect well-posedness for initial data in $\mathcal{Z}^{0}_{\rho}$, a problem that remains widely open, particularly in the non-decaying data setting.
The proof relies on a modified energy method [shatah, colliander-kdv] enabling us to control the derivative present in the nonlinearity of (gINLS), and to avoid the direct use of a standard fixed-point argument in the space $\mathcal{Z}^{2}_{\rho}$. We remark that in [Chen], using Kato’s iterative scheme similar to [deM-P], Chen claimed a well-posedness result in the same space $\mathcal{Z}^{2}_{\rho}$ for the defocusing Calogero-Moser DNLS equation (CM-defoc).
For the intermediate model (INLS), the analysis is more subtle because of the singular nature of the operator $T_{\delta}$. Although (INLS) is completely integrable for solutions in Sobolev spaces with vanishing boundary conditions, admits an inverse scattering transform and several Lax formulations [PG, CFL], these integrable structures do not provide well-defined conserved quantities in a setting with nonvanishing boundary conditions such as $\mathcal{Z}^{k}_{\rho}$. Indeed, if we consider the conservation laws at the $H^{n/2}$ level, derived via the Lax operator given in [CFL]:
$\int_{\mathbb{R}}\overline{u}\mathcal{L}_{u}^{n}u\,dx\quad(n=0,1,2,\dots),\qquad\mathcal{L}_{u}:\ f\in H^{1}(\mathbb{R})\mapsto-i\partial_{x}f+u\frac{1+iT_{\delta}}{2}(\overline{u}f)\in L^{2}(\mathbb{R}),$
then, after expansion and reduction of the number of $T_{\delta}$ operators using Cotlar’s identity (see [CFL, Eq. (2.12)]), the first five conservation laws read
@@
However, these integrals fail to be well-defined for $u\in\mathcal{Z}^{2}_{\rho}$, even after substituting $|u|^{2(n+1)}-\rho^{2(n+1)}$ for the last term $|u|^{2(n+1)}$ in $E_{n/2}(u)$. To overcome this, we construct conservation laws via the modified energy method, which allows us to establish global well-posedness for (gINLS). Our motivation for considering the more general model (gINLS), rather than (INLS), is to emphasize that the argument does not rely on the complete integrability of (INLS).
Theorem 1.2 (Global well-posedness and a priori bound).
Let $\delta>0$, $\rho>0$. For general $\alpha\neq 0$ and $\beta\geq 0$, the Cauchy problem associated with (gINLS) is globally well-posed in $\mathcal{Z}^{2}_{\rho}$. Moreover, for the integrable case (INLS), the following uniform bound holds:
$\sup_{t\in\mathbb{R}}\mathcal{E}^{2}_{\rho}(u(t))<\infty.$
In the general defocusing case $\alpha\neq 0$, $\beta\geq 0$, we still have a uniform bound for $\mathcal{E}^{1}_{\rho}(u(t))$ but only an exponential bound for $\|\partial_{x}^{2}u(t)\|_{L^{2}}$; see Theorem 4.1 in Section 4. The global theory is built on two new conservation laws compatible with nonvanishing boundary conditions, which allow us to control $\|u\|_{L^{\infty}}^{2},\|\partial_{x}u\|_{H^{1}}^{2},$ and $\||u|^{2}-\rho^{2}\|_{L^{2}}$ in time.
The paper is organized as follows. Section 2 collects preliminary estimates and properties of $T_{\delta}$ and $\mathcal{Z}^{k}_{\rho}$. Section 3 proves local well-posedness in $\mathcal{Z}^{2}_{\rho}$. Section 4 establishes the modified conservation laws and global well-posedness results.

### v2

# Global well-posedness for intermediate NLS with nonvanishing conditions at infinity

@@ 1. Introduction
We study the intermediate nonlinear Schrödinger equation in one space dimension,
@@
The equation (INLS) was introduced by Pelinovsky in [39] to describe the long-term dynamics of the envelope of weakly nonlinear internal waves in a stratified fluid of finite depth.
@@
Thus, as shown in [9, Theorem 1.3], when $\delta\to\infty$, (INLS) reduces to the defocusing Calogero-Moser derivative nonlinear Schrödinger equation
@@
originally derived on the real line in [39, 30, 19], and extended to the periodic setting in [34, 4, 6]. The equation (CM-defoc) has been deeply studied in recent years, where global well-posedness has been derived in the Hardy space $L^{2}_{+}(\mathbb{R})$ in [25] and in $L^{2}_{+}(\mathbb{T})$ in [4]. The focusing variant
@@
has also been shown to be globally well-posed in $L^{2}_{+}(\mathbb{R})$ [19, 25] for initial data $\|u(0)\|_{L^{2}}^{2}<2\pi$ and in $L^{2}_{+}(\mathbb{T})$ for small initial data [4], with the bound $2\pi$ known to be optimal thanks to the work of [22, 26, 27, 24]. Recently, Chapouto–Forlano–Laurens [9] proved that the well-posedness theory for (CM-foc) extends beyond the Hardy-space setting. They established local well-posedness in $H^{s}(\mathbb{R})$ for all $s>1/4$, and global well-posedness for initial data with small $L^{2}$-norm. Solitons for (CM-foc) were derived in [19, 30, 31, 36, 33] on the real line, and they display turbulent behavior [19]. On the torus, the set of special solutions, including traveling waves, is much wider than on $\mathbb{R}$ [6, 30, 36, 33]. We also mention that the zero dispersion limit on $\mathbb{R}$ was studied in [34, 5], and the inverse scattering theory for the Calogero-Moser model in the defocusing [35] and focusing [17, 10] cases has also been investigated.
@@
On the other hand, in the shallow-water limit $\delta\to 0$, after the rescaling $v(x,t)=\sqrt{\delta}u(x,t)$, (INLS) formally converges to the defocusing nonlinear Schrödinger equation
(NLS) $i\partial_{t}v-\partial_{x}^{2}v=-v|v|^{2}.$
@@
This equation and the focusing variant are famous integrable equations solved by the method of inverse scattering transformation [1, 43, 16, 29, 20]. Recently, Harrop-Griffiths–Killip–Vişan [21] proved that (NLS) and the focusing variant are globally well-posed in $H^{s}(\mathbb{R})$ for any s>1/2s>-1/2. This result is sharp in the sense that the norm inflation occurs in $H^{s}(\mathbb{R})$ for $s\leq-1/2$ (see, e.g., [12, 28]). See, e.g., [41, 21] for the development of well-posedness of nonlinear Schrödinger equations.
@@
A notable property of (INLS) is that it admits dark solitons [39, 37, 32]: for any $\rho>0$ and integer $N\geq 1$, there exists an $N$-soliton solution $Q$ which satisfies (INLS) on the whole time interval $\mathbb{R}$ and a nonvanishing boundary condition $|Q(x,t)|\to\rho$ as $|x|\to\infty$ for all $t\in\mathbb{R}$, while $\partial_{x}Q$ remains in $L^{2}(\mathbb{R})$. Thus, the existence of dark solitons naturally suggests that the appropriate functional framework for analyzing such solutions to (INLS) should involve nonvanishing boundary conditions. Accordingly, we introduce the following subsets of Zhidkov space:
(1.2) $\begin{gathered}\mathcal{Z}^{k}_{\rho}:=\{u\in L^{\infty}(\mathbb{R}):\partial_{x}u\in H^{k-1}(\mathbb{R}),\,|u|^{2}-\rho^{2}\in L^{2}(\mathbb{R})\}\qquad(\rho\geq 0),\\ \mathcal{Z}^{k}:=\bigcup_{\rho\geq 0}\mathcal{Z}^{k}_{\rho}=\{u\in L^{\infty}(\mathbb{R}):\partial_{x}u\in H^{k-1}(\mathbb{R}),\,|u|^{2}-\rho^{2}\in L^{2}(\mathbb{R})~\text{for some $\rho\geq 0$}\}.\end{gathered}$
@@
This framework naturally accommodates dark soliton solutions, which are localized deviation from a constant background. We will use the spaces $\mathcal{Z}^{k}$ only for positive integers $k$. Note that this is not a vector space but a complete metric space with the following distance function:
(1.3) $d^{k}(f,g):=\|\partial_{x}(f-g)\|_{H^{k-1}}+\|f-g\|_{L^{\infty}}+\big\|(|f|^{2}-\rho_{f}^{2})-(|g|^{2}-\rho_{g}^{2})\big\|_{L^{2}},$
where for $f\in\mathcal{Z}^{k}$ we denote by $\rho_{f}$ the unique non-negative number such that $|f|^{2}-\rho_{f}^{2}\in L^{2}(\mathbb{R})$. We also note that for each $\rho\geq 0$, $\mathcal{Z}^{k}_{\rho}$ is a closed subset of $\mathcal{Z}^{k}$, and that $H^{k}\hookrightarrow\mathcal{Z}^{k}_{0}$. To measure the size of functions in $\mathcal{Z}^{k}$, we introduce the functional
(1.4) $\mathcal{E}^{k}(u):=\|\partial_{x}u\|_{H^{k-1}}^{2}+\|u\|_{L^{\infty}}^{2}+\big\||u|^{2}-\rho_{u}^{2}\big\|_{L^{2}}.$
@@
Although the existence of dark solitons for (INLS) has been extensively studied in the literature [39, 37, 32, 30], and their asymptotic behavior has been investigated in [2], the problem of local well-posedness in a functional setting appropriate for these solitons remained open. In [15], local well-posedness222 Actually, he proved the local well-posedness of a family of nonlocal nonlinear Schrödinger equation [38] that includes also (INLS). was established in the Sobolev space $H^{s}(\mathbb{R})$ for $s\geq 1$. Later, Barros–de Moura–Santos have shown in [7] the local well-posedness of (INLS) for small initial data in the Besov space $B_{2}^{1/2,1}(\mathbb{R})$. More recently, Chapouto–Forlano–Laurens have proven in [9] that (INLS) is globally well-posed in $H^{s}(\mathbb{R})$ with $s>1/4$ for small $L^{2}$ initial data. However, this setting does not accommodate the nonvanishing boundary conditions intrinsic to dark solitons. The aim of the present work is therefore to address this gap by proving that (INLS) is globally well-posed in the space $\mathcal{Z}^{2}$. One could likely employ weighted Sobolev spaces to accommodate dark solitons; however, this approach lies beyond the scope of the present paper.
@@
Our first main result establishes local well-posedness in $\mathcal{Z}^{2}$ for (gINLS).
Theorem 1.1 (Local well-posedness in $\mathcal{Z}^{2}$).
Let $\delta>0$ and $\alpha,\beta\in\mathbb{R}$. For any $M_{0}>0$, there exists $T_{*}\in(0,1]$ satisfying
$T_{*}\geq c_{\alpha,\beta}\big[(1+\delta^{-1})(1+M_{0}^{4})\big]^{-1}$
such that for any initial datum $\phi\in\mathcal{Z}^{2}$ with $\mathcal{E}^{2}(\phi)\leq M_{0}^{2}$, the following hold:
  • There exists a unique solution $u\in C([0,T_{*}];\mathcal{Z}^{2})$ to (gINLS) with $u(0)=\phi$. Moreover, the solution map $\phi\mapsto u$ is continuous from $\{\phi\in\mathcal{Z}^{2}\colon\mathcal{E}^{2}(\phi)\leq M_{0}^{2}\}$ into $C([0,T_{*}];\mathcal{Z}^{2})$.
  • The solution preserves the background density: $u\in C([0,T_{*}];\mathcal{Z}^{2}_{\rho_{\phi}})$.
  • (Persistence of regularity) If $\phi\in\mathcal{Z}^{k}$ for some integer $k\geq 3$, then $u\in C([0,T_{*}];\mathcal{Z}^{k})$.
@@
maps a solution $u$ of (INLS) to another solution of (INLS), but with rescaled depth $\lambda\delta$. This formally indicates that $L^{2}(\mathbb{R})$ is the scaling-critical space. Consequently, one may expect well-posedness for initial data in $\mathcal{Z}^{0}$, a problem that remains widely open, particularly in the non-decaying data setting.
The proof relies on a modified energy method [40, 13] enabling us to control the derivative present in the nonlinearity of (gINLS), and to avoid the direct use of a standard fixed-point argument in the space $\mathcal{Z}^{2}$. We remark that in [11], using Kato’s iterative scheme similar to [14], Chen claimed a well-posedness result in the space $\mathcal{Z}^{2}_{\rho}$ with $\rho=1$ for the defocusing Calogero-Moser DNLS equation (CM-defoc). Observe that the existence time $T_{*}$ in Theorem 1.1 is uniformly bounded from below when $\delta\to\infty$. In fact, our proof is easily adapted to the limiting case $\delta=\infty$, yielding well-posedness in $\mathcal{Z}^{2}$ for the following generalized version of the Calogero-Moser DNLS equation:
(gCM) $i\partial_{t}u-\partial_{x}^{2}u=u(i\alpha-\beta H)\partial_{x}|u|^{2}\qquad\text{in $\mathbb{R}\times\mathbb{R}$}.$
Corollary 1.2 (Local well-posedness for (gCM)).
Let $\alpha,\beta\in\mathbb{R}$ and $M_{0}>0$. The same local well-posedness result as in Theorem 1.1 holds for (gCM), with $T_{*}\in(0,1]$ satisfying
$T_{*}\geq c_{\alpha,\beta}(1+M_{0}^{4})^{-1}.$
Beyond the independent well-posedness of both systems, the following result establishes the rigorous convergence of the (gINLS) flow to the (gCM) flow in the deep-water limit $\delta\to\infty$ for solutions with nonvanishing boundary conditions.
Corollary 1.3 (Convergence in the deep-water limit).
Let $\alpha,\beta\in\mathbb{R}$ be arbitrary parameters. Suppose that $\{\phi_{\delta}\}_{\delta\in[1,\infty]}\subset\mathcal{Z}^{2}$ is a family of initial data satisfying
$\lim_{\delta\to\infty}d^{2}(\phi_{\delta},\phi_{\infty})=0.$
Let $u_{\delta}$ for $\delta\in[1,\infty)$ and $u_{\infty}$ denote the unique maximal solutions in $C([0,T_{\delta,\max});\mathcal{Z}^{2})$ to the Cauchy problems for (gINLS) with $u_{\delta}(0)=\phi_{\delta}$ and (gCM) with $u_{\infty}(0)=\phi_{\infty}$, respectively, as constructed in Theorem 1.1 and Corollary 1.2. Then, the maximal existence time $T_{\delta,\max}$ satisfies the lower semicontinuity property:
$\liminf_{\delta\to\infty}T_{\delta,\max}\geq T_{\infty,\max}.$
Furthermore, for any $0<T<T_{\infty,\max}$, the solutions converge in the sense that
$\lim_{\delta\to\infty}\|d^{2}(u_{\delta},u_{\infty})\|_{L^{\infty}_{T}}=0.$
For the intermediate model (INLS), the analysis is more subtle than (NLS) and (CM-defoc) because of the singular nature of the operator $T_{\delta}$. Although (INLS) admits several Lax representations and infinite number of conservation laws [37, 9], these integrable structures do not provide well-defined conserved quantities in a setting with nonvanishing boundary conditions such as $\mathcal{Z}^{k}$. Indeed, if we consider the conservation laws at the $H^{n/2}$ level, derived via the Lax operator given in [9]:
(1.5) $\int_{\mathbb{R}}\overline{u}\mathcal{L}_{u}^{n}u\,dx\quad(n=0,1,2,\dots),\qquad\mathcal{L}_{u}:\ f\in H^{1}(\mathbb{R})\mapsto-i\partial_{x}f+u\frac{1+iT_{\delta}}{2}(\overline{u}f)\in L^{2}(\mathbb{R}),$
then, after expansion and reduction of the number of $T_{\delta}$ operators using Cotlar’s identity (see [9, Eq. (2.12)]), the first five conservation laws read
@@
However, these integrals fail to be well-defined for $u\in\mathcal{Z}^{2}$ with $\rho_{u}>0$, even after substituting $|u|^{2(n+1)}-\rho_{u}^{2(n+1)}$ for the last term $|u|^{2(n+1)}$ in $E_{n/2}(u)$. To overcome this, we construct conservation laws via the modified energy method, which allows us to establish global well-posedness for (gINLS). Our motivation for considering the more general model (gINLS), rather than (INLS), is to emphasize that the argument does not rely on the complete integrability of (INLS).
Theorem 1.4 (Global well-posedness and a priori bounds).
Let $\delta>0$. For general $\alpha\neq 0$ and $\beta\geq 0$, the Cauchy problem is globally well-posed in
$\mathcal{Z}^{2}_{*}:=\bigcup_{\rho>0}\mathcal{Z}^{2}_{\rho}$
for both (gINLS) and (gCM). Furthermore, the energy functional satisfies a uniform bound in the integrable–defocusing cases:
$\sup_{t\in\mathbb{R}}\mathcal{E}^{2}\big(u(t)\big)<\infty,$
which holds for (INLS) as well as for (CM-defoc).
Remark 1.2.
In the general defocusing case $\alpha\neq 0$, $\beta\geq 0$, we still have a uniform bound for $\mathcal{E}^{1}(u(t))$ but only an exponential bound for $\|\partial_{x}^{2}u(t)\|_{L^{2}}$; see Theorem 4.1 in Section 4. The global theory is built on two new conservation laws compatible with nonvanishing boundary conditions, which allow us to control $\|u\|_{L^{\infty}}^{2},\|\partial_{x}u\|_{H^{1}}^{2},$ and $\||u|^{2}-\rho_{u}^{2}\|_{L^{2}}$ in time.
The paper is organized as follows. Section 2 collects preliminary estimates and properties of $T_{\delta}$ and $\mathcal{Z}^{k}$. Section 3 proves local well-posedness in $\mathcal{Z}^{2}$. Section 4 establishes the modified conservation laws and global well-posedness in $\mathcal{Z}^{2}_{*}$.